OpenStudy (anonymous):

what does sin2x x sin2x =? (Remember it's not the same as 2sinx x 2sinx)

OpenStudy (agent0smith):

\[\large \sin (2x) \sin (2x) = \sin^2 (2 x)\]

OpenStudy (agent0smith):

You can also change it into some other crap if you prefer, using double angle formulas: http://home.windstream.net/okrebs/page103.html

Parth (parthkohli):

And by the way,\[\large \sin(2x)\sin(2x) \]\[\large = 4\sin^2(x)\cos^2(x) \]\[\large = \sin^2(2x)\]\[\large = 4(\cos^2(x) - \cos^4(x))\]\[\large = 4(\sin^2(x) - \sin^4(x)) \]

Parth (parthkohli):

Oh, and\[\large = 1 - \cos^2(2x)\]